# Chiral susceptibility and axial U(1) anomaly near the (pseudo-)critical temperature

### v1

@@ 4 Summary
We thank L. Glozman and Y. Sumino for useful discussions. For the numerical simulation we use the QCD software packages Grid [16, 18] for configuration generations and Bridge++ [17, 19] for measurements. Numerical simulations were performed on Oakforest-PACS and Wisteria/BDEC-01 at JCAHPC under a support of the HPCI System Research Projects (Project IDs: hp170061, hp180061, hp190090, and hp200086, hp210104, hp220093, hp230070) as well as Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba (Project IDs: xg17i032, xg18i023 and wo22i038) and Fugaku computer provided by the RIKEN Center for Computational Science under a support of the HPCI System Research Projects (Project IDs: hp200130, hp210165, hp210231, hp220279, hp230323) We used Japan Lattice Data Grid (JLDG) [20] for storing a part of the numerical data generated for this work. This work is supported in part by the Japanese Grant-in-Aid for Scientific Research (No. JP26247043, JP18H01216, JP18H04484, JP18H05236, JP22H01219) and by Joint Institute for Computational Fundamental Science (JICFuS).

### v2

@@ 4 Summary
We thank L. Glozman and Y. Sumino for useful discussions. For the numerical simulation we have used the QCD software packages Grid [16, 18] for configuration generations and Bridge++ [17, 19] for measurements. Numerical simulations were performed on Oakforest-PACS and Wisteria/BDEC-01 at JCAHPC under a support of the HPCI System Research Projects (Project IDs: hp170061, hp180061, hp190090, and hp200086, hp210104, hp220093, hp230070) as well as Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba (Project IDs: xg17i032, xg18i023 and wo22i038) and Fugaku computer provided by the RIKEN Center for Computational Science under a support of the HPCI System Research Projects (Project IDs: hp200130, hp210165, hp210231, hp220279, hp230323). This work was (partly) achieved through the use of SQUID at the Cybermedia Center, Osaka University. We used Japan Lattice Data Grid (JLDG) [20] for storing a part of the numerical data generated for this work. This work is supported in part by the Japanese Grant-in-Aid for Scientific Research (No. JP26247043, JP18H01216, JP18H04484, JP18H05236, JP22H01219) and by Joint Institute for Computational Fundamental Science (JICFuS).